A package with mass M is unloaded from a Truck with an inclined ramp. The ramp has rollers that eliminates friction and the truck unnotes the package from an initial height H. The ramp is inclined at an angle B. Find the algebraic expression in terms of this quantities for the work done on the package during its trip down the ramp.
1
Expert's answer
2019-09-19T09:46:40-0400

The force will be:


"F=mg"

The displacement will be:


"l=\frac{h}{\sin{B}}"

The work done on the package during its trip down the ramp:


"W=Fl\cos{(90\degree-B)}=Fl\sin{B}=mg\frac{h}{\sin{B}}\sin{B}"

"W=mgh"